OpenStudy (anonymous):

if a convex polyhedron has 6 vertices and 8 faces then it has _____ edges?

OpenStudy (anonymous):

using the Euler's Formula F + V - E = 2, can be used for all polyhedral. F = number of faces, V = number of vertices, E = number of edges so; 8 + 6 - E = 2 14 - e = 2 -E = 2 - 12 -E = -12 E = 12

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np :)